OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. ****Write the solution as an ordered pair. I know it looks like the same question as the previous one I posted, but this one is different than the other one from the bottom. To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. ****Write the solution as an ordered pair. I know it looks like the same question as the previous one I posted, but this one is different than the other one from the bottom. @Mathematics

OpenStudy (anonymous):

we can write the amount a person must pay as a function in terms of the flat rate+session fee*number of sessions (n) \[f(n)=m+f*n\]

OpenStudy (anonymous):

when he had attended n=5 sessions we have 25=m+5f and when he had attended n=10 sessions we have 40=m+10f subtract the second equation from the first and we get 15=5f 3=f substituting that in to either equation we get 25=m+5(3) 10=m or 40=m+10(3) 10=m so our solution is m=10, f=3, or (m,f)=(10,3)

OpenStudy (across):

Perhaps I'm missing something, but why do you add the membership fee to the second session?

OpenStudy (anonymous):

he paid 40 dollars in all, meaning membership + session fees

OpenStudy (across):

In all here meaning something akin to a cumulative payment?

OpenStudy (across):

If that's the case, wouldn't then 'm' be better replaced by 25? I don't know, I'm a little confused.

OpenStudy (anonymous):

in all meaning m+nf meaning all payments towards it. it is the equivalent of saying a different person had been there 10 times and payed 40 dollars, but m shouldn't be 25 because the membership fee is not constant b/c otherwise he would have payed 25 dollars at both times